# Preserving Task-Relevant Information Under Linear Concept Removal

### v1

@@ 4.2 Language modeling
We focus on two language modeling tasks. First, we use a dataset from Limisiewicz et al. (2024), which we refer to as the profession dataset. Inspired by Bolukbasi et al. (2016), this dataset contains prompt templates containing professions, which need to be finished by the LM (e.g.,‘the plumber wanted that’). Each profession has a stereotype score $z_{\mathrm{stereo}}$. This indicates how strongly a profession is connected with the male gender through stereotypical cues (e.g., plumber has a high stereotype score, while nurse has a low one). Each profession also has a factual score $y_{\mathrm{fact}}$, which indicates how strong a profession is connected to the male gender through factual information (e.g., waiter has a high factual score, but waitress has a low one).
@@
Set-up of the experiments: for the profession dataset, our goal is to create an LM that does not rely on stereotypical cues, but on factual information. We estimate the extent to which the LM $\mathcal{M}$ relies on stereotypical cues or factual information as follows. Let $t_{\mathrm{he}}$/$t_{\mathrm{she}}$ be the tokens for "he"/"she". Let $\bm{t}_{i}$ denote tokens for a prompt $i$, and $p_{\mathcal{M}}(t_{\mathrm{he}}|\bm{t}_{i})$ the probability assigned by a model of the "he" token conditional on the prompt $\bm{t}_{i}$. We measure the log-odds ratio between the probability of the next token being ‘he’ or ‘she’ as
@@
Results: for the experiment on the profession dataset, the results are shown in Table 1. We report the exponent of the coefficients in Equation 7, as this tells us how more likely the ‘he’ token becomes relative to the ‘she’ token after a one-unit increase in either the stereotypical or factual score. After applying any of the three projections, the extent to which the model relies on stereotypical information is greatly reduced, per the reduction in $\exp(\hat{\beta}_{\mathrm{stereo}})$. The extent to which the model relies on factual information after a projection is greatly reduced when applying SAL or LEACE, whereas it is increased or preserved after applying SPLICE.
@@
For the experiment on the Winobias dataset, the results are shown in Figure 3. For two out of three Llama models, SPLICE improves coreference accuracy more than the other projections. In particular, it strongly increases the accuracy for anti-stereotypical prompts.
@@ 4.3 Application to image data
In Figure 4 we illustrate the effect of each projection on the raw pixels, for several images. SPLICE accentuates parts of the image that are useful for distinguishing images with and without glasses. For instance, it tends to make the areas around the eyes lighter when someone does not wear glasses, and darken when they do. This shows that SPLICE mitigates the damage to the “glasses” features exposed to a linear classifier, despite of the high correlation. We illustrate that this holds on average across the whole dataset in Appendix B.3.
@@ 5 Discussion and Limitations
Theoretically, we show that the range of the projection—particularly, whether or not it includes $\bm{\Sigma}_{\bm{x},\bm{y}}$—can matter only if one does not re-fit the last linear layer, or refit it with a regularization term. However, we lack a theory that would explain under which conditions it does matter, and when it is expected to outperform LEACE. Future work should study whether preserving main-task covariance is optimal under some settings. In this paper, we focus on preserving that quantity as it is intuitively related to the desideratum of main-task performance; however, we note that the SPLICE objective can easily be modified to preserve any direction that is not identical to the covariance with the protected attribute.
In addition, a limitation of the SPLICE objective is that it prioritizes preserving $\bm{\Sigma}_{\bm{x},\bm{y}}$ over minimizing $\mathbb{E}\left[\big{\|}\mathbf{P}\bm{x}-\bm{x}\big{\|}^{2}_{\mathbf{M}}\right]$. This can potentially lead to distortive changes to the embeddings. We investigate this limitation in Appendix B.4. As the SPLICE objective sometimes fails when intervening in middle layers, we aim to study the adaption of the SPLICE objective for preserving the directions in the representation space that are being used by an LM in some middle layer.
@@ 6 Conclusion
We introduce SPLICE, a method that generalizes previous concept-erasure methods by provably removing the ability to linearly predict sensitive information, while maintaining the covariance between the representations and another main-task label. Our analysis pins the problem down to a pair of geometric constraints—placing $\mathrm{colsp}(\bm{\Sigma}_{\bm{x},\bm{z}})$ in the kernel of the projection while forcing $\mathrm{colsp}(\bm{\Sigma}_{\bm{x},\bm{y}})$ to lie in its range—and proves that the oblique projector of Theorem 1 is the unique minimum-distortion solution under these constraints. Experimentally, SPLICE tends to better preserve average and worst-group accuracy on the Bias in Bios and Multilingual Text Detoxification tasks when the task–concept correlation is high. In a language-modeling setting, we are able to influence stereotypical bias while preserving factual gender information; and in most models, it is better in preserving LM’s ability to perform co-reference after debiasing in the Winobias dataset. Future work should formalize whether maintaining the covariance with the main-task translates into main-task loss guarantees and develop variants over the SPLICE objective that impose weaker distortion when applied to earlier hidden layers.
@@ A.1 Proof of theorem 1
Let $m=\mathrm{rank}(\bm{\Sigma}_{\bm{x},\bm{x}}),$ with $1\leq m\leq d.$ Let $\mathcal{W}=\mathrm{colsp}(\mathbf{W})$ be the subspace in $\mathbb{R}^{d}$ in which $\bm{x}$ has non-zero variance. Without loss of generality, we will define a basis for $\bm{x}$ where the first $m$ coordinates of $\bm{x}$ lie in $\mathcal{W}$. The last $l=d-m$ coordinates lie in its orthogonal complement $\mathcal{W}^{\perp}$. Our $\bm{x}$ can now be written as
@@
where $p=i-l+1.$ We use the indexation via $p$ because we use the rows $p\in\{1,...,l\}$ of the matrices $\mathbf{C}$ and $\mathbf{D}$ respectively to represent $(\mathbf{P}\bm{x})_{i}$. The objective in equation A.1 for $i\in\{l,...,d\}$ and corresponding $p$ can be written as
@@
where in the second equality we used that $\check{x}_{i}=0$ almost surely and in the final equality we used that $\mathbb{E}[\tilde{x}_{i}]=0.$ We note that the values for $\mathrm{D}_{p,j}$ do not matter for the objective. Furthere, because $\mathrm{Cov}(\tilde{\bm{x}},\tilde{\bm{x}})$ is p.s.d., we can achieve the minimum of equation 17 by setting $\mathbf{C}=\mathbf{0}.$ For simplicity, we also set $\mathbf{D}=\mathbf{0}.$ This then also trivially satisfies the kernel and range constraints for the components $i\in\{l,...,d\}.$
@@
$\mathcal{C}_{1}$ =$=\left\{M\in\mathbb{R}^{m\times m}\mid\tilde{\mathbf{W}}M\tilde{\mathbf{W}}\bm{\Sigma}_{\tilde{\bm{x}},\bm{z}}=\mathbf{0}_{m}\right\},$ (23b)
$\mathcal{C}_{2}$ =$=\left\{M\in\mathbb{R}^{m\times m}\mid\tilde{\mathbf{W}}M\tilde{\mathbf{W}}\bm{\Sigma}_{\tilde{\bm{x}},\bm{y}}=\tilde{\mathbf{W}}\bm{\Sigma}_{\tilde{\bm{x}},\bm{y}}\right\}.$ (23c)
@@
This corresponds to $\mathbf{P}^{\star}_{\mathrm{SPLICE}}$ in equation 4 and concludes the proof of Theorem 1.

### v2

@@ 4.2 Language modeling
We focus on two language modelling tasks using the Llama series of models (Touvron et al., 2023; Grattafiori et al., 2024). First, we use a dataset from Limisiewicz et al. (2024), which we refer to as the profession dataset. Inspired by Bolukbasi et al. (2016), this dataset contains prompt templates containing professions, which need to be finished by the LM (e.g.,‘the plumber wanted that’). Each profession has a stereotype score $z_{\mathrm{stereo}}$. This indicates how strongly a profession is connected with the male gender through stereotypical cues (e.g., plumber has a high stereotype score, while nurse has a low one). Each profession also has a factual score $y_{\mathrm{fact}}$, which indicates how strong a profession is connected to the male gender through factual information (e.g., waiter has a high factual score, but waitress has a low one).
@@
Set-up of the experiments: for the profession dataset, our goal is to create an LM that does not rely on stereotypical cues, but on factual information. We estimate the extent to which the LM $\mathcal{M}$ relies on stereotypical cues or factual information as follows. Let $t_{\mathrm{he}}$$t_{\mathrm{she}}$ be the tokens for "he"/"she". Let $\bm{t}_{i}$ denote tokens for a prompt $i$, and $p_{\mathcal{M}}(t_{\mathrm{he}}|\bm{t}_{i})$ the probability assigned by a model of the "he" token conditional on the prompt $\bm{t}_{i}$. We measure the log-odds ratio between the probability of the next token being ‘he’ or ‘she’ as
@@
Results: for the experiment on the profession dataset, the results are shown in Table 1. We report the exponent of the coefficients in Equation 7, as this tells us how more likely the ‘he’ token becomes relative to the ‘she’ token after a one-unit increase in either the stereotypical or factual score. After applying any of the three projections, the extent to which the model relies on stereotypical information is greatly reduced, per the reduction in $\exp(\hat{\beta}_{\mathrm{stereo}})$. The extent to which the model relies on factual information after a projection is greatly reduced when applying SAL or LEACE, whereas it is increased or preserved after applying SPLINCE.
@@
For the experiment on the Winobias dataset, the results are shown in Figure 3. For two out of three Llama models, SPLINCE improves coreference accuracy more than the other projections. In particular, it strongly increases the accuracy for anti-stereotypical prompts. In Appendix B.6 report results for additional LM’s outside of the Llama series.
@@ 4.3 Application to image data
In Figure 4 we illustrate the effect of each projection on the raw pixels, for several images. SPLINCE accentuates parts of the image that are useful for distinguishing images with and without glasses. For instance, it tends to make the areas around the eyes lighter when someone does not wear glasses, and darken when they do. This shows that SPLINCE mitigates the damage to the “glasses” features exposed to a linear classifier, despite of the high correlation. We illustrate that this holds on average across the whole dataset in Appendix B.3.
In Appendix B.5 we include additional results CelebA dataset as introduced here, as well as the Waterbirds dataset Sagawa et al. (2020). SPLINCE performs relatively worse for these vision classification tasks than the NLP classification tasks in 4.1. This gap in performance between the vision and NLP classification tasks is an interesting direction for future research.
@@ 5 Discussion and Limitations
Theoretically, we show that the range of the projection—particularly, whether or not it includes $\bm{\Sigma}_{\bm{x},\bm{y}}$—can matter only if one does not re-fit the last linear layer, or refit it with a regularization term. However, we lack a theoretical understanding that would explain under which conditions it does matter, and when it is expected to outperform LEACE. In Appendix A.3 we provide an initial investigation into this question, showing SPLINCE is guaranteed to outperform LEACE when we freeze the last layer and the embeddings are whitened. Future work should study whether preserving main-task covariance is optimal under more general settings. In this paper, we focus on preserving that quantity as it is intuitively related to main-task performance; however, we note that the SPLINCE objective can be modified to preserve any direction that is not identical to the covariance with the protected attribute.
In addition, a limitation of the SPLINCE objective is that it prioritizes preserving $\bm{\Sigma}_{\bm{x},\bm{y}}$ over minimizing $\mathbb{E}\left[\big\|\mathbf{P}\bm{x}-\bm{x}\big\|^{2}_{\mathbf{M}}\right]$. This can potentially lead to distortive changes to the embeddings. We investigate this limitation in Appendix B.4. As the SPLINCE objective sometimes fails when intervening in middle layers, we aim to study the adaption of the SPLINCE objective for preserving the directions in the representation space that are being used by an LM in some middle layer.
While we did not investigate a multi-modal setting (e.g. CLIP, Radford et al. (2021)), one potential limitation of SPLINCE is that covariance subspaces might not be aligned across modalities.
Finally, see Appendix D for a discussion on ethical considerations when applying SPLINCE.
@@ 6 Conclusion
We introduce SPLINCE, a method that generalizes previous concept-erasure methods by provably removing the ability to linearly predict sensitive information, while maintaining the covariance between the representations and another main-task label. Our analysis pins the problem down to a pair of geometric constraints—placing $\mathrm{colsp}(\bm{\Sigma}_{\bm{x},\bm{z}})$ in the kernel of the projection while forcing $\mathrm{colsp}(\bm{\Sigma}_{\bm{x},\bm{y}})$ to lie in its range—and proves that the oblique projector of Theorem 1 is the unique minimum-distortion solution under these constraints. Experimentally, SPLINCE tends to better preserve average and worst-group accuracy on the Bias in Bios and Multilingual Text Detoxification tasks when the task–concept correlation is high. In a language-modeling setting, we are able to influence stereotypical bias while preserving factual gender information; and in most models, it is better in preserving LM’s ability to perform co-reference after debiasing in the Winobias dataset. Future work should formalize whether maintaining the covariance with the main-task translates into main-task loss guarantees and develop variants over the SPLINCE objective that impose weaker distortion when applied to earlier hidden layer, or performs better for vision datasets.
@@ A.1 Proof of theorem 1
Let $m=\mathrm{rank}(\bm{\Sigma}_{\bm{x},\bm{x}}),$ with $1\leq m\leq d.$ Let $\mathcal{W}=\mathrm{colsp}(\mathbf{W})$ be the subspace in $\mathbb{R}^{d}$ in which $\bm{x}$ has non-zero variance. Without loss of generality, we will define a basis for $\bm{x}$ where the first $m$ coordinates of $\bm{x}$ lie in $\mathcal{W}$. The last l=dml=d-m coordinates lie in its orthogonal complement $\mathcal{W}^{\perp}$. Our $\bm{x}$ can now be written as
@@
where p=il+1.p=i-l+1. We use the indexation via $p$ because we use the rows $p\in\{1,...,l\}$ of the matrices $\mathbf{C}$ and $\mathbf{D}$ respectively to represent $(\mathbf{P}\bm{x})_{i}$. The objective in equation A.1 for $i\in\{l,...,d\}$ and corresponding $p$ can be written as
@@
[…]
@@
$\mathcal{C}_{1}$ $=\left\{M\in\mathbb{R}^{m\times m}\mid\tilde{\mathbf{W}}M\tilde{\mathbf{W}}\bm{\Sigma}_{\tilde{\bm{x}},\bm{z}}=\mathbf{0}_{m}\right\},$ (23b)
$\mathcal{C}_{2}$ $=\left\{M\in\mathbb{R}^{m\times m}\mid\tilde{\mathbf{W}}M\tilde{\mathbf{W}}\bm{\Sigma}_{\tilde{\bm{x}},\bm{y}}=\tilde{\mathbf{W}}\bm{\Sigma}_{\tilde{\bm{x}},\bm{y}}\right\}.$ (23c)
@@
This corresponds to $\mathbf{P}^{\star}_{\mathrm{SPLINCE}}$ in equation 4 and concludes the proof of Theorem 1.
@@
A.3 Excess risk of SPLINCE vs. LEACE without re-training the last layer
In this section, we provide two theorems that show conditions under which SPLINCE is guaranteed to outperform LEACE when the last layer is frozen (e.g. not re-trained, contrasting the set-up of Theorem 3.2). In Theorem A.3 we prove that for a linear regression with whitened data from any distribution, SPLINCE does not degrade the performance of a frozen last layer. In contrast, applying LEACE may lead to an increase in the loss. In Theorem A.3 we prove a similar statement, but for the case where the embeddings $\bm{x}$ follow a standard multivariate Gaussian distribution.
We emphasize that both theorems are limited in their applicability, as we generally would not expect last-layer embeddings to be whitened or follow a standard multivariate Gaussian distribution. Further work is required to understand more general conditions when SPLINCE might outperform LEACE.
{restatable}
[Excess-risk of leace and splince in a regression setting] theoremExcessRiskThmRegression Let $\mathrm{\mathbf{X}}\in\mathbb{R}^{n\times d}$ be whitened data with $\bm{\Sigma}_{\bm{x},\bm{x}}=\bm{I}_{d}$ and $\mathbb{E}[\bm{x}]=\mathbf{0}_{d}$. Assume the response model $\mathrm{\mathbf{Y}}=\mathrm{\mathbf{X}}\bm{\beta}+\bm{\epsilon}$, where $\mathbb{E}[\bm{\epsilon}]=\mathbf{0}_{n}$, $\operatorname{Var}(\bm{\epsilon})=\sigma^{2}\bm{I}_{n}$ and $\mathbb{E}[\bm{\epsilon}^{\top}\mathrm{\mathbf{X}}]=\mathbf{0}_{d}^{\mathrm{T}}$.
Let $\mathbf{Q}_{\textsc{leace}}^{\top}:=\bm{I}_{d}-\mathbf{P}_{\bm{z}}$, where $\mathbf{P}_{\bm{z}}$ projects onto the subspace spanned by a concept variable $\bm{z}$; let $\mathbf{Q}_{\textsc{splince}}^{\top}:=\mathbf{P}$ be a projection satisfying $\bm{\Sigma}_{\bm{x},\bm{y}}\in\operatorname{Range}(\mathbf{P})$. Then:
  1. 1.
    The excess risk of leace is greater than or equal to zero, $\Delta R(\mathbf{Q}_{\textsc{leace}})\geq 0.$
  2. 2.
    The excess risk of splince is zero, $\Delta R(\mathbf{Q}_{\textsc{splince}})=0$.
Proof.
For any square matrix $\mathbf{Q}$ define the risk
$R(\mathbf{Q}):=\mathbb{E}\bigl[(\mathrm{\mathbf{Y}}-\mathrm{\mathbf{X}}\mathbf{Q}^{\top}\bm{\beta})^{2}\bigr],$
and the excess risk
$\Delta R(\mathbf{Q}):=R(\mathbf{Q})-R(\bm{I}_{d}).$
Using $\mathrm{\mathbf{Y}}=\mathrm{\mathbf{X}}\bm{\beta}+\bm{\epsilon}$ we have
$\mathrm{\mathbf{Y}}-\mathrm{\mathbf{X}}\mathbf{Q}^{\top}\bm{\beta}$ $=\mathrm{\mathbf{X}}\bm{\beta}+\bm{\epsilon}-\mathrm{\mathbf{X}}\mathbf{Q}^{\top}\bm{\beta}$
$=\bm{\epsilon}+\mathrm{\mathbf{X}}(\bm{I}_{d}-\mathbf{Q}^{\top})\bm{\beta}.$ (1)
Hence
$R(\mathbf{Q})$ $=\mathbb{E}\Bigl[(\bm{\epsilon}+\mathrm{\mathbf{X}}(\bm{I}_{d}-\mathbf{Q}^{\top})\bm{\beta})^{\!\top}(\bm{\epsilon}+\mathrm{\mathbf{X}}(\bm{I}_{d}-\mathbf{Q}^{\top})\bm{\beta})\Bigr]$
$=\underbrace{\mathbb{E}[\bm{\epsilon}^{\top}\bm{\epsilon}]}_{\sigma^{2}}+2\,\underbrace{\mathbb{E}\bigl[\bm{\epsilon}^{\top}\mathrm{\mathbf{X}}(\bm{I}_{d}-\mathbf{Q}^{\top})\bm{\beta}\bigr]}_{0}+\mathbb{E}\Bigl[\bm{\beta}^{\top}(\bm{I}_{d}-\mathbf{Q})\mathrm{\mathbf{X}}^{\top}\mathrm{\mathbf{X}}(\bm{I}_{d}-\mathbf{Q}^{\top})\bm{\beta}\Bigr]$
$=\sigma^{2}+\bm{\beta}^{\top}(\bm{I}_{d}-\mathbf{Q})\bm{\Sigma}_{\bm{x},\bm{x}}(\bm{I}_{d}-\mathbf{Q}^{\top})\bm{\beta}$
$=\sigma^{2}+\|(\bm{I}_{d}-\mathbf{Q}^{\top})\bm{\beta}\|_{2}^{2},$ (2)
because $\bm{\Sigma}_{\bm{x},\bm{x}}=\bm{I}_{d}$.
Subtracting $R(\bm{I}_{d})=\sigma^{2}$ from (2) yields the expression
$\boxed{\;\Delta R(\mathbf{Q})=\|(\bm{I}-\mathbf{Q}^{\top})\bm{\beta}\|_{2}^{2}\;}.$
Take $\mathbf{Q}=\mathbf{Q}_{\textsc{leace}}=\bm{I}_{d}-\mathbf{P}_{\bm{z}}$. Since $\mathbf{P}_{\bm{z}}$ is a projector, $(\bm{I}_{d}-\mathbf{P}_{\bm{z}})^{\top}=\bm{I}_{d}-\mathbf{P}_{\bm{z}}$; thus
$\Delta R(\mathbf{Q}_{\textsc{leace}})=\|\mathbf{P}_{\bm{z}}\bm{\beta}\|_{2}^{2}.$
Unless $\mathbf{P}_{\bm{z}}\bm{\beta}=\mathbf{0}_{d}$ (the degenerate case where $\bm{\beta}$ lies fully outside the concept subspace), this quantity is strictly positive.
Let $\mathbf{Q}=\mathbf{Q}_{\textsc{splince}}=\mathbf{P}^{\top}$. By assumption $\bm{\Sigma}_{\bm{x},\bm{y}}\in\operatorname{Range}(\mathbf{P})$, and under whitening $\bm{\beta}=\bm{\Sigma}_{xy}$. Hence $\mathbf{P}\bm{\beta}=\bm{\beta}$ and $(\bm{I}_{d}-\mathbf{P}^{\top})\bm{\beta}=(\bm{I}_{d}-\mathbf{P})\bm{\beta}=\mathbf{0}_{d}.$ Applying the boxed identity, $\Delta R(\mathbf{Q}_{\textsc{splince}})=0$. ∎
{restatable}
[Excess-risk bound of leace and splince in a logistic-regression setting] theoremExcessRiskThmLogistic Let $\mathrm{\mathbf{X}}\in\mathbb{R}^{n\times d}$ contain i.i.d. rows $\bm{x}\sim\mathcal{N}\!\bigl(\mathbf{0}_{d},\bm{\Sigma}_{\bm{x},\bm{x}}\bigr)$. Assume the conditional model
$\mathbb{P}(y=1\mid\bm{x})=g\!\bigl(\mathbb{\bm{x}}^{\!\top}\bm{\beta}\bigr),\qquad g(s)=\tfrac{1}{1+e^{-s}},$
with true parameter $\bm{\beta}\in\mathbb{R}^{d}$.
Define $\mathbf{Q}_{\textsc{leace}}^{\top}:=\bm{I}_{d}-\mathbf{P}_{\bm{z}}$, where $\mathbf{P}_{\bm{z}}$ projects onto the subspace spanned by a concept variable $\bm{z}$; define $\mathbf{Q}_{\textsc{splince}}^{\top}:=\mathbf{P}$, where $\mathbf{P}$ is an orthogonal projector satisfying $\bm{\Sigma}_{\bm{x},\bm{y}}\in\operatorname{Range}(\mathbf{P})$.
Let $\Delta_{\!\ell}(\mathbf{Q}):=\mathbb{E}[\ell(\mathbb{\bm{x}}^{\!\top}\mathbf{Q}^{\!\top}\bm{\beta},y)-\ell(\mathbb{\bm{x}}^{\!\top}\bm{\beta},y)]$ denote the (population) excess logistic risk, with $\ell(s,y)=\log(1+e^{s})-ys$.
Then:
  1. 1.
    $0<\Delta_{\!\ell}\!\bigl(\mathbf{Q}_{\textsc{leace}}\bigr)\;\leq\;\frac{1}{8}\,\bm{\beta}^{\!\top}\mathbf{P}_{\bm{z}}\,\bm{\Sigma}_{\bm{x},\bm{x}}\,\mathbf{P}_{\bm{z}}\bm{\beta}.$
    In the whitened case $\bm{\Sigma}_{\bm{x},\bm{x}}=\bm{I}_{d}$ this reduces to $\Delta_{\!\ell}(\mathbf{Q}_{\textsc{leace}})\leq\tfrac{1}{8}\lVert\mathbf{P}_{\bm{z}}\bm{\beta}\rVert_{2}^{2}.$
  2. 2.
    (SPLINCE) $\Delta_{\!\ell}\!\bigl(\mathbf{Q}_{\textsc{splince}}\bigr)=0.$
Proof.
Define
$s:=\bm{x}^{\!\top}\bm{\beta},\qquad\hat{s}:=\bm{x}^{\!\top}\mathbf{Q}^{\!\top}\bm{\beta},\qquad\delta(x):=\hat{s}-s=\bm{x}^{\!\top}(\mathbf{Q}^{\!\top}-\bm{I}_{d})\bm{\beta}=\bm{x}^{\!\top}(\bm{I}_{d}-\mathbf{Q}^{\!\top})(-\bm{\beta}).$
Three facts about the logistic loss.
  1. Fact 1.
    $\ell^{\prime}(s,y)=\sigma(s)-y$.
  2. Fact 2.
    $\ell^{\prime\prime}(s,y)=\sigma(s)\bigl(1-\sigma(s)\bigr)\leq\frac{1}{4}.$
  3. Fact 3.
    By the mean–value theorem, for some $\theta=\theta(x,y)\in(0,1)$,
    $\ell(\hat{s},y)=\ell(s,y)+\delta\,\ell^{\prime}(s,y)+\frac{\delta^{2}}{2}\,\ell^{\prime\prime}\!\bigl(s-\theta\delta,y\bigr).$
Taking expectations and using Fact 3,
$\Delta_{\!\ell}(Q)$ $=\mathbb{E}\bigl[\ell(\hat{s},y)-\ell(s,y)\bigr]$ (35)
$=\mathbb{E}\!\bigl[\delta\,\ell^{\prime}(s,y)\bigr]+\frac{1}{2}\,\mathbb{E}\!\bigl[\delta^{2}\ell^{\prime\prime}\!\bigl(s-\theta\delta,y\bigr)\bigr].$ (36)
Using $\mathbb{E}[y\mid x]=\sigma(s)$ (by model assumption),
$\mathbb{E}\!\bigl[\delta\,\ell^{\prime}(s,y)\bigr]$ $=\mathbb{E}_{x}\!\Bigl[\delta(x)\,\underbrace{\mathbb{E}[\,\sigma(s)-y\mid x]}_{=0}\Bigr]=0.$ (37)
Fact 2 says $\ell^{\prime\prime}\leq\frac{1}{4}$; hence
$\frac{1}{2}\,\mathbb{E}\bigl[\delta^{2}\ell^{\prime\prime}(\cdot)\bigr]\;\leq\;\frac{1}{8}\,\mathbb{E}\!\bigl[\delta^{2}\bigr].$ (38)
Because $\bm{x}\sim\mathcal{N}(\mathbf{0}_{d},\bm{\Sigma}_{\bm{x},\bm{x}})$,
$\mathbb{E}\bigl[\delta^{2}\bigr]$ $=\bm{\beta}^{\!\top}(\bm{I}_{d}-\mathbf{Q})\,\bm{\Sigma}_{\bm{x},\bm{x}}\,(\bm{I}_{d}-\mathbf{Q}^{\!\top})\bm{\beta}.$ (39)
Thus
 $\boxed{\;0\leq\Delta_{\!\ell}(\mathbf{Q})\leq\frac{1}{8}\,\bm{\beta}^{\!\top}(\bm{I}_{d}-\mathbf{Q})\bm{\Sigma}_{\bm{x},\bm{x}}(\bm{I}_{d}-\mathbf{Q}^{\!\top})\bm{\beta}\;}.$ (40)
For Gaussian $x$ and differentiable $g$, Stein’s lemma states $\mathbb{E}[g(x)\,x]=\bm{\Sigma}_{\bm{x},\bm{x}}\,\mathbb{E}[\nabla g(x)]$. Taking $g(x)=\sigma(\bm{x}^{\!\top}\bm{\beta})$ (a scalar function) gives
$\bm{\Sigma}_{\bm{x},\bm{y}}$ $=\mathbb{E}[\bm{x}y]=\mathbb{E}[\bm{x}\,\sigma(s)]=\bm{\Sigma}_{\bm{x},\bm{x}}\,\mathbb{E}\!\bigl[\sigma^{\prime}(s)\,\bm{\beta}\bigr]$ (41)
$=\bm{\Sigma}_{\bm{x},\bm{x}}\,\bm{\beta}\,\underbrace{\mathbb{E}\!\bigl[\sigma(s)(1-\sigma(s))\bigr]}_{=:C}.$ (42)
Because $C>0$, we can solve for the true parameter:
$\bm{\beta}=\frac{1}{C}\,\bm{\Sigma}_{\bm{x},\bm{x}}^{-1}\bm{\Sigma}_{\bm{x},\bm{y}}.$ (43)
Substituting yields for any $\mathbf{Q}$,
$\Delta_{\!\ell}(\mathbf{Q})\leq\frac{1}{8C^{2}}\,\bm{\Sigma}_{\bm{x},\bm{y}}^{\!\top}\bm{\Sigma}_{\bm{x},\bm{x}}^{-1}(\bm{I}_{d}-\mathbf{Q})\bm{\Sigma}_{\bm{x},\bm{x}}(\bm{I}_{d}-\mathbf{Q}^{\!\top})\bm{\Sigma}_{\bm{x},\bm{x}}^{-1}\bm{\Sigma}_{\bm{x},\bm{y}}.$ (44)
Take $\mathbf{Q}=\mathbf{Q}_{\textsc{leace}}=\bm{I}_{d}-\mathbf{P}_{\bm{z}}$.  Since $(\bm{I}_{d}-\mathbf{Q})=\mathbf{P}_{\bm{z}}$ is a projector,
$\Delta_{\!\ell}(\mathbf{Q}_{\textsc{leace}})\leq\frac{1}{8C^{2}}\,\bigl\lVert\mathbf{P}_{\bm{z}}\bm{\Sigma}_{\bm{x},\bm{y}}\bigr\rVert^{2}.$
Unless $\mathbf{P}_{\bm{z}}\bm{\Sigma}_{\bm{x},\bm{y}}=\mathbf{0}_{d}$ (degenerate), the RHS is strictly positive, proving the first half of the theorem.
Let $\mathbf{Q}=\mathbf{Q}_{\textsc{splince}}=\mathbf{P}^{\top}$ where $\mathbf{P}\bm{\Sigma}_{\bm{x},\bm{y}}=\bm{\Sigma}_{\bm{x},\bm{y}}$. Then $(\bm{I}_{d}-\mathbf{Q}^{\!\top})\bm{\Sigma}_{\bm{x},\bm{y}}=(\bm{I}_{d}-\mathbf{P})\bm{\Sigma}_{\bm{x},\bm{y}}=0$, so the right-hand side of equation 44 is zero. Because $\Delta_{\!\ell}(\mathbf{Q})\geq 0$ by definition, $\Delta_{\!\ell}(\mathbf{Q}_{\textsc{splince}})=0$. ∎